OpenStudy (anonymous):

One microgram is 10^6 times smaller than 1 gram. If the sugar refinery makes 1014 grams of sugar a month, what is the weight of that sugar in micrograms? 10^84 10^8 10^20 10^-8

OpenStudy (anonymous):

@robtobey @TheSmartOne @iambatman @StudyGurl14 @satellite73 @Donblue22 @ganeshie8 @jigglypuff314 @KendrickLamar2014 @Nnesha

OpenStudy (studygurl14):

Do you mean \(10^{14}\), not 1014?

OpenStudy (studygurl14):

If so...then just multiply \(10^{14}\) and \(10^6\) together.

OpenStudy (kendricklamar2014):

10^6 = ? @judtpleasegahd do u now what that is <-----

OpenStudy (studygurl14):

Rule of exponents: \(\large x^mx^n=x^{m(n)}\)

OpenStudy (kendricklamar2014):

10^6 = 1000000

OpenStudy (studygurl14):

You don't even have to convert to standard form @KendrickLamar2014 . It's easier just to deal with the exponents.

OpenStudy (anonymous):

@StudyGurl14 no how its written is how its suppose to be

OpenStudy (kendricklamar2014):

So 1 Micro gram is 1,000 times smaller than 1 Gram

OpenStudy (studygurl14):

Um...okay....then I'm assuming they are rounded 1014 to 1000....which is written as \(10^3\) In that case, multiply \(10^3\) and \(10^6\) together.

OpenStudy (studygurl14):

Which is not an answer choice, so I don't think it is 1014. I think it is \(10^{14}\)

OpenStudy (anonymous):

@StudyGurl14 thats not an answer choice

OpenStudy (kendricklamar2014):

10^14 = 100,000,000,000,000

OpenStudy (studygurl14):

What's not...?

OpenStudy (studygurl14):

Just in case you're mistaken, I'm not saying \(10^{14}\) is the answer... @TheSmartOne can you help me here?

OpenStudy (studygurl14):

I'm obviously not explaining it well enough... @TheSmartOne

OpenStudy (kendricklamar2014):

10^14 * 10^6 = 100,000,000,000,000,000,000

OpenStudy (anonymous):

@StudyGurl14 my mistake you right it is 10^14

OpenStudy (studygurl14):

Okay, so... \(10^{14}\times10^6\tt = the~answer\)

OpenStudy (anonymous):

A?

OpenStudy (studygurl14):

correct

OpenStudy (anonymous):

it was C :(

OpenStudy (anonymous):

but thanks

OpenStudy (studygurl14):

Oh, oops..sorry I meant to put \(x^mx^n=x^{m+n}\)

OpenStudy (kendricklamar2014):

Sorry, I took so long, it was C, becuz 10^14 + 10^6 = 10^20

OpenStudy (studygurl14):

I'm so...so...SO sorry @judtpleasegahd

OpenStudy (anonymous):

its ok can you help me with my other question

OpenStudy (studygurl14):

Sure. I promise not to mess up this time. :)

OpenStudy (anonymous):

http://openstudy.com/study#/updates/546020e1e4b0c2835b864be0 lol its ok @StudyGurl14